OpenStudy (shamallamadingdong):

Any help is much appreciated At the curb, a ramp is 11 inches off the ground. The other end of the ramp rests on the street 55 inches from the curb. Write a linear equation in slope-intercept form that relates the height y of the ramp to the distance x from the curb.

OpenStudy (shamallamadingdong):

@paki any good at math

OpenStudy (shamallamadingdong):

y=1/5x+55 y=-1/5x+55 y=-1/5x+11 y=5x+11

OpenStudy (shamallamadingdong):

@pooja195 @Elsa213 ;-;

pooja195 (pooja195):

Since slope is rise/run, the slope has to be 11 (the height, or rise of the ramp) over 55 ( the length, or run of the ramp) Also, since the problem states the highest part of the ramp first, from left the right the ramp is going down, so the slope is negative. 11//55 reduces to 1/5, and the slope must be negative, and the y intercept is always the starting point there for its 11 (where the ramp starts) So whats the answer?

OpenStudy (shamallamadingdong):

ilyu pooja answer is c

pooja195 (pooja195):

Correct :) ilyut

OpenStudy (shamallamadingdong):

thank u for not blocking me

OpenStudy (shamallamadingdong):

<3

pooja195 (pooja195):

xD <3